# LEADRE: Multi-Faceted Knowledge Enhanced LLM Empowered Display Advertisement Recommender System

### v1

@@ 6. Conclusion
In this work, we present the first industrial application of generative retrieval in display advertising. To address the challenges of ”How to capture user commercial intent”, ”How to bridge the gap between LLMs and ad-Specific knowledge”, and ”How to efficiently deploy LLMs”, we introduce a novel LLM-based framework called Multi-Faceted Knowledge Enhanced LLM Empowered Display ADvertisement REcommender system (LEADRE). LEADRE comprises three core modules: the Intent-Aware Prompt Engineering, the Advertising-Specific Knowledge Alignment, and the Efficient System Deployment. To evaluate the effectiveness of LEADRE, we conducted extensive experiments, including both offline and online evaluations. The results demonstrate its ability to generate more accurate and diverse ads. Moreover, online A/B test revealed that LEADRE resulted in a 1.57% increase on Tencent WeChat Channels, as well as a 1.17% increase on Tencent WeChat Moments in GMV for serviced users.
@@
Appendix A Preliminary
A.1. Ads Indexing
The RQ-VAE takes the ad’s textual embedding $\mathbf{x}_{a}$ as input and uses an encoder-quantization-decoder mechanism to generate a list of discrete IDs. In the encoding phase, the textual embedding $\mathbf{x}_{a}$ is encoded by an encoder, $f_{En}(\cdot;\theta_{En})$, to obtain a hidden embedding, formulated as $\hat{\mathbf{z}}_{a}=f_{En}(\mathbf{x}_{a};\theta_{En})$, where $\hat{\mathbf{z}}_{a}\in\mathbb{R}^{d_{RQ}}$ denotes the hidden embedding, and $d_{RQ}$ is the hidden embedding’s dimension. During the quantization phase, we apply $M$ layers of residual vector quantization. For each layer $l$, we maintain a codebook $\mathcal{C}^{l}=\{\mathbf{e}_{k}^{l}\}_{k=1}^{K}$, where $\mathbf{e}_{k}^{l}\in\mathbb{R}^{d_{RQ}}$ and $K$ denotes the number of codes in the codebook. The quantization process is defined as follows:
$c^{l}_{a}=\text{argmin}_{k}||\mathbf{r}^{l}_{a}-\mathbf{e}_{k}^{l}||_{2}^{2},\quad\mathbf{r}^{l+1}_{a}=\mathbf{r}^{l}_{a}-\mathbf{e}_{c^{l}_{a}}^{l},\quad\mathbf{z}_{a}=\sum_{l=1}^{M}\mathbf{e}_{c^{l}_{a}}^{l}$
$\mathcal{L}_{recons}=\sum_{a}||\mathbf{x}_{a}-\hat{\mathbf{x}}_{a}||_{2}^{2},$
$\mathcal{L}_{quant}=\sum_{a}\sum_{l=1}^{M}\left(||\text{sg}[\mathbf{r}^{l}_{a}]-\mathbf{e}_{c^{l}_{a}}^{l}||_{2}^{2}+\beta_{quant}||\mathbf{r}^{l}_{a}-\text{sg}[\mathbf{e}_{c^{l}_{a}}^{l}]||_{2}^{2}\right),$
This process ensures that ads are represented by discrete S-IDs, which allows the LLMs to generate valid ads from the predefined ads set during constrained decoding.
A.2. Trie-Tree Construction
We employ a trie-tree to represent all ads, where each path from the root node to a leaf node corresponds to the S-IDs of a specific ad. The construction of the trie-tree follows the Trie-Tree Construction Algorithm. The algorithm initializes the trie-tree with a head node and processes each ad $a$ in the set $\mathcal{A}$ by sequentially examining its S-ID list $\mathbf{c}_{a}$ from the root.
At each step, the algorithm checks whether the current S-ID $c_{a}^{l}$ already exists as a child of the current node. If the S-ID is not present, a new node is created and added as a child, labeled with the current S-ID. The algorithm then updates the current node to the child node corresponding to the processed S-ID. After processing all S-IDs of an ad, the current node is marked as the end of the ad, indicating that a complete ad sequence has been stored. This process is repeated for all ads in the set $\mathcal{A}$. Once all ads have been processed, the algorithm returns the constructed trie-tree, which efficiently organizes the ads in a hierarchical structure. This structure allows for quick retrieval of ads based on their S-IDs and supports efficient prefix-based searches.
A.3. Constrained Decoding
The Constrained Decoding algorithm leverages a trie-tree and beam search to ensure that the ads generated by LLMs conform to a predefined set of valid ads. The trie-tree, encompassing all possible ad S-IDs, constrains the LLM’s output to valid S-IDs.
The algorithm initializes a generated ad beam set comprising $B$ lists, each starting with the head node of the trie-tree. For each layer of the trie-tree, corresponding to the position in the S-ID list, the algorithm retrieves valid next S-IDs based on the current node in the trie-tree. The LLM offers a probability distribution over feasible next S-IDs, constrained by the trie-tree structure. For each valid next S-ID, the algorithm appends the S-ID to the current candidate list, computes the score using the LLM’s output probabilities, and adds the updated list to a new beam set. After processing all candidate lists in the current beam set, the top $B$ lists are selected based on their scores, and the process continues to the next layer of the trie-tree. This process is repeated until either all layers of the trie have been processed or the maximum S-IDs length $(M+1)$ is reached. The algorithm returns the final beam set, which contains the generated ads that are valid according to the trie-tree and have the highest probabilities based on the LLM’s predictions.
By combining beam search with trie-tree constraints, the algorithm ensures that the generated ads are both high-probability lists according to the LLM and valid ads from the predefined ads set.

### v3

@@ 6. Conclusion
In this work, we present the first industrial application of generative retrieval in display advertising. To address the challenges of ”How to capture user commercial intent”, ”How to bridge the gap between LLMs and ad-Specific knowledge”, and ”How to efficiently deploy LLMs”, we introduce a novel LLM-based framework called Multi-Faceted Knowledge Enhanced LLM Empowered Display ADvertisement REcommender system (LEADRE). LEADRE comprises three core modules: the Intent-Aware Prompt Engineering, the Advertising-Specific Knowledge Alignment, and the Latency-Aware Model Deployment. To evaluate the effectiveness of LEADRE, we conducted extensive experiments, including both offline and online evaluations. The results demonstrate its ability to generate more accurate and diverse ads. Moreover, online A/B test revealed that LEADRE resulted in a 1.57% increase on Tencent WeChat Channels, as well as a 1.17% increase on Tencent WeChat Moments in GMV for serviced users.
@@
Acknowledgments
This work was supported by National Natural Science Foundation of China (62436010, 62172421) and the Tencent Innovation Fund. Biao Qin is the first corresponding author.